# Text Indexing and Pattern Matching with Ephemeral Edits

### v1

@@ 3.3 Edit Operation
Let us start by characterizing the type of occurrences of $P$ in $T^{i}$. We will use this to argue about the correctness of our algorithm. We have the following occurrence types:
@@
For insertions or substitutions, all the above occurrence types are relevant, whereas for deletions only Types 1, 2, and 5 are relevant. Inspect Fig. 5 for an illustration. It is readily verifiable that these are all possible occurrence types.
@@
For starters, let us assume that $\epsilon=1$ and suppose that an edit operation happens at position $p$ of $T$. We first consider the case of a letter deletion; i.e., we are asked to delete letter $T[p]$. We find the occurrences of $P$ occurring in $T^{i}=T[0\mathinner{.\,.}p)T[p+1\mathinner{.\,.}n)$ as follows. We first find the occurrences of $P$ in $T[0\mathinner{.\,.}p)$ (Type 1) and in $T[p+1\mathinner{.\,.}n)$ (Type 2) by using the standard procedure for 3-sided queries: we retrieve the SA interval of $P$ in $T$ (we have this from the preprocessing part of $P$) and use RmQ and RMQ queries recursively on the interval until the answer of a query exceeds the bounds [0,pm+1][0,p-m+1] (for RmQs) and $[p+1,n)$ (for RMQs). This takes $\mathcal{O}(1+\textsf{Occ})$ time because every query takes $\mathcal{O}(1)$ time [BF00] and we stop after two unsuccessful queries (one RmQ and one RMQ) or if the intervals are empty. We still need to find the occurrences spanning $T[\mathinner{.\,.}p-1]\cdot T[p+1\mathinner{.\,.}]$ (Type 5). We make the following observation. If we had the longest prefix of $P$ occurring as a suffix of $T[0\mathinner{.\,.}p)$ and the longest suffix of $P$ occurring as a prefix of $T[p+1\mathinner{.\,.}n)$, then we could use a single prefix-suffix query, which takes $\mathcal{O}(1)$ time (Theorem 4), to report all the occurrences of $P$ spanning $T[\mathinner{.\,.}p-1]\cdot T[p+1\mathinner{.\,.}]$. For locating the longest such suffix of $P$ (finding the prefix is symmetric by reversing $T$ and $P$ and so we omit the details), we observe the following.
@@
Suppose that we had the suffix tree of T#PT\#P, $\notin\Sigma$, and that we had marked the nodes of the suffix tree representing the suffixes of $P$. The longest suffix of $P$ starting at position $p+1$ of $T$ is the deepest marked ancestor of the leaf representing $T[p+1\mathinner{.\,.}n)$ in the suffix tree of T#PT\#P.
@@
We now consider the case of inserting the letter $c$ after $T[p]$. We need to find the occurrences of $P$ spanning $T[\mathinner{.\,.}p]cT[p+1\mathinner{.\,.}]$ (Type 5) using a prefix-suffix query. However, for insertions, things are a bit more complicated as we need to locate the longest suffix of $P$ that occurs as a prefix of $cT[p+1\mathinner{.\,.}n)$. We make the following observation for locating the longest such suffix of $P$ (finding the prefix is analogous to deletion, and so we omit it).
@@
Suppose that we had the suffix tree of T#PT\#P, $\notin\Sigma$, and that we had marked the nodes of the suffix tree representing the suffixes of $P$. The longest suffix $P[i-1]P[i\mathinner{.\,.}m)$ of $P$ starting at position $p+1$ of $T^{i}$, for some $i$, is such that the string $P[i\mathinner{.\,.}m)$ is represented by the deepest marked ancestor of the leaf representing $T[p+1\mathinner{.\,.}n)$ in the suffix tree of T#PT\#P and $P[i-1]=T^{i}[p+1]$.
@@
We find the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$ as follows. When $p+1$ is given, we find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}n)$ among all suffixes of $T$. We then search for $r$ in the predecessor structure for group $c$ to infer $s,e]_{i}$, such that $r\in[s,e]_{i}$, if it exists. This gives us the longest suffix $cP[i\mathinner{.\,.}m)$ of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$. Recall that to find the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ we do the same work as for the deletion operation on the reverse of $P$ and $T$. After finding the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ and the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$, we trigger one prefix-suffix query. The total time to process the insertion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\log\log n+\textsf{Occ})$, where the $\log\log n$ term is due to predecessor search. The algorithm is correct by 2.
@@
We have that $T^{i}=T[0\mathinner{.\,.}q)T[p+1\mathinner{.\,.}n)$. For the occurrences of $P$ in $T[0\mathinner{.\,.}q)$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}n)$ (Type 2) we use RMQs similar to $\epsilon=1$. For the occurrences of $P$ with a starting position in $[0,q)$ and an ending position in $q,|T^{i}|)$ (Type 5), we find the longest suffix of $P$ which is a prefix of $T[p+1\mathinner{.\,.}n)$ and the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}q)$ again similar to $\epsilon=1$. The total time for processing the deletion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\log\log n+\textsf{Occ})$.
@@
We have that $T^{i}=T[0\mathinner{.\,.}p]\cdot S\cdot T[p+1\mathinner{.\,.}n)$. For the occurrences of $P$ in $T[0\mathinner{.\,.}p)$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}n)$ (Type 2) we use RMQs similar to $\epsilon=1$. For the occurrences of $P$ with a starting position in $[0,p]$ and an ending position after the edit (Type 5), we find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}n)$ in the predecessor structure coming from $\textsf{TREE}_{S}(P)$. This gives us the longest suffix of $P$ occurring as a prefix of $S\cdot T[p+1\mathinner{.\,.}n)$. We also find the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$, just like for $\epsilon=1$, and trigger one prefix-suffix query. We are not done yet, as we may have an occurrence of $P$ with a starting position in $[p+1,p+|S|]$ and an ending position in $p+|S|+1,|T^{i}|)$ (Type 4). We find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}n)$ in the predecessor structure derived from $\textsf{TREE}(P)$. This gives us the longest suffix of $P$ occurring as a prefix of $T[p+1\mathinner{.\,.}n)$. We also find the longest prefix of $P$ that is a suffix of $S$ naively in $\mathcal{O}(\epsilon^{2})=\mathcal{O}(1)$ time, and trigger another prefix-suffix query. (Type 3 is symmetric to Type 4.) Further note that Type 6 occurrences can be found in $\epsilon=\mathcal{O}(1)$ time using any linear-time pattern matching algorithm [KJP77]. The total time to process the insertion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\log\log n+\textsf{Occ})$.
@@
Finally, note that substitutions work the same as insertions, but with a different offset. Namely, the only change is that when fragment $T[p\mathinner{.\,.}p+|S|-1]$ is substituted by string $S$, we find the lex-rank $r=\textsf{iSA}[p+|S|]$ of $T[p+|S|\mathinner{.\,.}n)$ in the predecessor structure derived from $\textsf{TREE}_{S}(P)$; and then we find the lex-rank $r=\textsf{iSA}[p+|S|]$ of $T[p+|S|\mathinner{.\,.}n)$ in the predecessor structure coming from $\textsf{TREE}(P)$.
After reporting the occurrences of $P$ in $T^{i}$, we do nothing (to revert the edit); i.e., we proceed to the next edit or preprocess a new pattern. We have arrived at Theorem 1.
@@ 4 Pattern Matching with Ephemeral Block Deletions
We design a data structure for pattern matching with ephemeral block deletions. First, we preprocess two strings $T$ and $P$, each of length at most $n$, over alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$. Then, we allow any ephemeral sequence of arbitrarily-long block deletions in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time. This improves upon Theorem 1 for this setting.
@@ 4.1 Preprocessing
We then employ 1 (inspect Fig. 6). We construct the suffix tree $\textsf{ST}(T\#P)$ of T#PT\#P, $\notin\Sigma$, and mark the nodes of the suffix tree that represent the suffixes of $P$. The longest suffix of $P$ occurring as a prefix of $T[p\mathinner{.\,.}|T|)$, for all $p$, is the deepest marked ancestor of the leaf representing $T[p\mathinner{.\,.}n)$ in $\textsf{ST}(T\#P)$. Using a bottom-up traversal on the suffix tree and a stack to maintain the currently active marked nodes, we store a pointer from every leaf representing a suffix of $T$ to its deepest marked ancestor (i.e., the longest suffix of $P$ represented by an ancestor of the leaf). By reversing $T$ and $P$, we also store the longest prefix of $P$ which is a suffix of the prefix $T[0\mathinner{.\,.}p]$ of $T$, for all $p$. This preprocessing takes $\mathcal{O}(n)$ time and space [FFM00].
@@
Figure 6: The suffix tree of $T\#P=\texttt{bababbbababb}\#\texttt{ababab}$. The label of edges leading to leaf nodes is truncated to avoid clattering the figure. The nodes representing suffixes of $P$ are colored red. The longest suffix of $P$ occurring as a prefix of $T[p\mathinner{.\,.}|T|)$ for $p=7$ is the node decorated by $15$.
@@ 4.2 Block Deletion
Again, it is readily verifiable that only Type 1, 2, and 5 occurrences are relevant. Assume that fragment $T[q\mathinner{.\,.}p]$ is deleted from $T$. In particular, we have $T^{i}=T[0\mathinner{.\,.}q)T[p+1\mathinner{.\,.}|T|)$. In addition, assume that $|T^{i}|\geq|P|$. (If this is not the case, no occurrence of $P$ in $T^{i}$ is to be reported.) For the occurrences of $P$ in $T[0\mathinner{.\,.}q)$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}|T|)$ (Type 2) we use RMQs, similar to Section 3. Namely, each occurrence is reported in $\mathcal{O}(1)$ time. For the occurrences of $P$ with a starting position in $[0,q)$ and an ending position in $q,|T^{i}|)$ (Type 5), we retrieve the longest suffix of $P$ that is a prefix of $T[p+1\mathinner{.\,.}|T|)$ and the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}q)$, and trigger a prefix-suffix query. The total time to process deletion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\textsf{Occ}_{S})$. After reporting the occurrences of $P$ in $T^{i}$, we do nothing (to revert the edit). The correctness follows from 1. We have arrived at Theorem 2.
@@
5 Pattern Matching with Ephemeral Edits
We design a data structure for pattern matching with ephemeral edit operations. First, we preprocess two strings $T$ and $P$ each of length at most $n$, over alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$. Then, we allow any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time. This improves upon Theorem 1 for this setting.
@@ 5.1 Preprocessing
We start by constructing the data structure of Theorem 4 for prefix-suffix queries over $P$. This takes $\mathcal{O}(|P|)$ time and space. We also construct the String Matching Automaton (SMA) of the reverse of $P$. The SMA of a string $X$, denoted by $\textsf{SMA}(X)$, is the minimal deterministic automaton that accepts the language $\Sigma^{*}X$, where $X\in\Sigma^{*}$. It has $|X|+1=\mathcal{O}(|X|)$ states, one for every prefix of $X$, with $\varepsilon$ being the initial state and $X$ being the accepting state. The crucial property of SMA is that only the transitions that do not lead to the initial state are stored. There are at most $2|X|=\mathcal{O}(|X|)$ of them, regardless of the size of the alphabet [CHL07]. The $\textsf{SMA}(X)$ can be constructed in $\mathcal{O}(|X|)$ time [CHL07]; its transitions can be accessed in $\mathcal{O}(1)$ time if stored using perfect hashing [BCF+23]. This preprocessing takes $\mathcal{O}(|P|)$ time and space.
Deletions are handled by Theorem 2; we shift our focus to preprocessing for insertions. We employ 2 (inspect Fig. 7). We construct the suffix tree $\textsf{ST}(T\#P)$ of T#PT\#P, $\notin\Sigma$, and mark the nodes of the suffix tree that represent the suffixes of $P$. Every marked node representing $P[i\mathinner{.\,.}|P|)$ has a pointer to the state of SMA representing the reverse of $P[i\mathinner{.\,.}|P|)$. Using a bottom-up traversal on the suffix tree and a stack to maintain the currently active marked nodes, we store a pointer from every leaf representing a suffix of $T$ to its deepest marked ancestor (i.e., the longest suffix of $P$ represented by an ancestor of the leaf). We also store the longest prefix of $P$ which is a suffix of the prefix $T[0\mathinner{.\,.}p]$ of $T$, for all $p$ (see Section 4.1). This preprocessing takes $\mathcal{O}(n)$ time and space.
@@
Figure 7: On the left is the suffix tree of $T\#P=\texttt{bababbbababb}\#\texttt{ababab}$. The label of edges leading to leaf nodes is truncated to avoid clattering the figure. The nodes representing suffixes of $P$ are colored red. On the top right is the SMA of the reverse of $P$. Consider the query $p=4$ and $c=\texttt{a}$. The pointer from the leaf node labeled $p+1=5$ of the suffix tree takes us to its deepest colored ancestor labeled 1818; the corresponding state in the SMA is 1. From there, if we follow the a transition, it takes us forward to state 2, meaning that the longest suffix of $P$ preceded by $c=\texttt{a}$, represented by an ancestor of the leaf node labeled 55, is b, which yields ab.

### v2

@@ 3.3 Edit Operation
Let us start by classifying the occurrences of $P$ in $T^{i}$. We will use this classification to argue about the correctness of our algorithm. We have the following occurrence types:
@@
For insertions or substitutions, all the above occurrence types are relevant, whereas for deletions only Types 1, 2, and 5 are relevant. Inspect Figure 6 for an illustration. It is readily verifiable that these are all possible occurrence types.
@@
For starters, let us assume that $\epsilon=1$ and suppose that an edit operation happens at position $p$ of $T$. We first consider the case of a letter deletion; i.e., we are asked to delete letter $T[p]$. We find the occurrences of $P$ occurring in $T^{i}=T[0\mathinner{.\,.}p)T[p+1\mathinner{.\,.}n)$ as follows. We first find the occurrences of $P$ in $T[0\mathinner{.\,.}p)$ (Type 1) and in $T[p+1\mathinner{.\,.}n)$ (Type 2) by using the standard procedure for 3-sided queries: we retrieve the SA interval of $P$ in $T$ (we have this from the preprocessing part of $P$) and use RmQ and RMQ queries recursively on the interval until the answer of a query exceeds the bounds [0,pm][0,p-m] (for RmQs) and [p+1,nm][p+1,n-m] (for RMQs). This takes $\mathcal{O}(1+\textsf{Occ})$ time because every query takes $\mathcal{O}(1)$ time [BF00] and we stop after two unsuccessful queries (one RmQ and one RMQ) or if the intervals are empty.
We still need to find the occurrences spanning $T[\mathinner{.\,.}p-1]\cdot T[p+1\mathinner{.\,.}]$ (Type 5). We make the following observation. If we had the longest prefix of $P$ occurring as a suffix of $T[0\mathinner{.\,.}p)$ and the longest suffix of $P$ occurring as a prefix of $T[p+1\mathinner{.\,.}n)$, then we could use a single prefix-suffix query, which takes $\mathcal{O}(1)$ time (Theorem 4), to report all the occurrences of $P$ spanning $T[\mathinner{.\,.}p-1]\cdot T[p+1\mathinner{.\,.}]$. For locating the longest such suffix of $P$ (finding the prefix is symmetric by reversing $T$ and $P$ and so we omit the details), we observe the following.
@@
Suppose that we had the suffix tree of T#PT\#P, $\notin\Sigma$, and that we had marked the nodes of the suffix tree representing the suffixes of $P$. The longest suffix of $P$ starting at position $p+1$ of $T$ is the deepest marked ancestor of the node representing $T[p+1\mathinner{.\,.}n)$ in the suffix tree of T#PT\#P.
@@
We now consider the case of inserting the letter $c$ after $T[p]$. We need to find the occurrences of $P$ spanning $T[\mathinner{.\,.}p]cT[p+1\mathinner{.\,.}]$ (Type 5) using a prefix-suffix query. However, for insertions, things are a bit more complicated: we need to locate the longest suffix of $P$ that occurs as a prefix of $cT[p+1\mathinner{.\,.}n)$. We make the following observation for locating the longest such suffix of $P$ (finding the prefix is analogous to deletion, and so we omit it).
@@
Suppose that we had the suffix tree of T#PT\#P, $\notin\Sigma$, and that we had marked the nodes of the suffix tree representing the suffixes of $P$. The longest suffix $P[i-1]P[i\mathinner{.\,.}m)$ of $P$ starting at position $p+1$ of $T^{i}$, for some $i$, is such that the string $P[i\mathinner{.\,.}m)$ is represented by the deepest marked ancestor of the node representing $T[p+1\mathinner{.\,.}n)$ in the suffix tree of T#PT\#P and $P[i-1]=T^{i}[p+1]$.
@@
We find the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$ as follows. When $p+1$ is given, we find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}n)$ among all suffixes of $T$. We then search for $r$ in the predecessor structure for group $c$ to infer $s,e]_{i}$, such that $r\in[s,e]_{i}$, if it exists. This gives us the longest suffix $cP[i\mathinner{.\,.}m)$ of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$. Recall that to find the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ we do the same work as for the deletion operation. After finding the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$ and the longest suffix of $P$ occurring as a prefix of $cT[p+1\mathinner{.\,.}n)$, we trigger one prefix-suffix query. The total time to process the insertion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\log\log n+\textsf{Occ})$, where the $\log\log n$ term is due to predecessor search. The algorithm is correct by 2.
@@
We have that $T^{i}=T[0\mathinner{.\,.}q)T[p+1\mathinner{.\,.}n)$. For the occurrences of $P$ in $T[0\mathinner{.\,.}q)$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}n)$ (Type 2) we use RMQs similar to $\epsilon=1$. For the remaining occurrences of $P$ (Type 5), we find the longest suffix of $P$ which is a prefix of $T[p+1\mathinner{.\,.}n)$ and the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}q)$ again similar to $\epsilon=1$. The total time for processing the deletion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\log\log n+\textsf{Occ})$.
@@
We have that $T^{i}=T[0\mathinner{.\,.}p]\cdot S\cdot T[p+1\mathinner{.\,.}n)$. For the occurrences of $P$ in $T[0\mathinner{.\,.}p]$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}n)$ (Type 2) we use RMQs similar to $\epsilon=1$. For the occurrences of $P$ with a starting position in $[0,p]$ and an ending position after the edit (Type 5), we find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}n)$ in the predecessor structure coming from $\textsf{TREE}_{S}(P)$. This gives us the longest suffix of $P$ occurring as a prefix of $S\cdot T[p+1\mathinner{.\,.}n)$. We also find the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}p]$, just like for $\epsilon=1$, and trigger one prefix-suffix query. We are not done yet, as we may have an occurrence of $P$ with a starting position in $[p+1,p+|S|]$ and an ending position in $p+|S|+1,|T^{i}|)$ (Type 4). We find the lex-rank $r=\textsf{iSA}[p+1]$ of $T[p+1\mathinner{.\,.}n)$ in the predecessor structure derived from $\textsf{TREE}(P)$. This gives us the longest suffix of $P$ occurring as a prefix of $T[p+1\mathinner{.\,.}n)$. We also find the longest prefix of $P$ that is a suffix of $S$ naively in $\mathcal{O}(\epsilon^{2})=\mathcal{O}(1)$ time, and trigger another prefix-suffix query. (Type 3 is symmetric to Type 4.) Further note that Type 6 occurrences can be found in $\epsilon=\mathcal{O}(1)$ time using any linear-time pattern matching algorithm [KJP77]. The total time to process the insertion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\log\log n+\textsf{Occ})$.
@@
Finally, note that substitutions work the same as insertions, but with a different offset: the only change is that when fragment $T[p\mathinner{.\,.}p+|S|-1]$ is substituted by string $S$, we find the lex-rank $r=\textsf{iSA}[p+|S|]$ of $T[p+|S|\mathinner{.\,.}n)$ in the predecessor structure derived from $\textsf{TREE}_{S}(P)$; and then we find the lex-rank $r=\textsf{iSA}[p+|S|]$ of $T[p+|S|\mathinner{.\,.}n)$ in the predecessor structure coming from $\textsf{TREE}(P)$. After reporting the occurrences of $P$ in $T^{i}$, we do nothing (to revert the edit); i.e., we proceed to the next edit or preprocess a new pattern. We have arrived at Theorem 1.
@@ 4 Pattern Matching with Ephemeral Block Deletions
In this section, we prove Theorem 2. We design a data structure for pattern matching with ephemeral block deletions. First, we preprocess two strings $T$ and $P$, each of length at most $n$, over alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$. Then, we allow any ephemeral sequence of arbitrarily-long block deletions in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time. This improves upon Theorem 1 for this setting.
@@ 4.1 Preprocessing
We then employ 1 (inspect Figure 7). We construct the suffix tree $\textsf{ST}(T\#P)$ of T#PT\#P, $\notin\Sigma$, and mark the nodes of the suffix tree that represent the suffixes of $P$. The longest suffix of $P$ occurring as a prefix of $T[p\mathinner{.\,.}|T|)$, for all $p$, is the deepest marked ancestor of the node representing $T[p\mathinner{.\,.}|T|)$ in $\textsf{ST}(T\#P)$. Using a bottom-up traversal on the suffix tree and a stack to maintain the currently active marked nodes, we store a pointer from every node $v$ representing a suffix of $T$ to its deepest marked ancestor (i.e., the longest suffix of $P$ represented by an ancestor of node $v$). By reversing $T$ and $P$, we also store the longest prefix of $P$ that is a suffix of the prefix $T[0\mathinner{.\,.}p]$ of $T$, for all $p$. This preprocessing takes $\mathcal{O}(n)$ time and space [FFM00].
@@
Figure 7: The suffix tree of $T\#P=\texttt{bababbbababb}\#\texttt{ababab}$. The label of edges leading to leaf nodes is truncated to avoid cluttering the figure. The nodes representing suffixes of $P$ are colored red. The longest suffix of $P$ occurring as a prefix of $T[p\mathinner{.\,.}|T|)$ for $p=7$ is the node associated with $15$.
@@ 4.2 Block Deletion
Again, it is readily verifiable that only Type 1, 2, and 5 occurrences are relevant. Assume that fragment $T[q\mathinner{.\,.}p]$ is deleted from $T$. In particular, we have $T^{i}=T[0\mathinner{.\,.}q)T[p+1\mathinner{.\,.}|T|)$. In addition, assume that $|T^{i}|\geq|P|$. (If this is not the case, no occurrence of $P$ in $T^{i}$ is to be reported.) For the occurrences of $P$ in $T[0\mathinner{.\,.}q)$ (Type 1) we use RmQs and for the occurrences of $P$ in $T[p+1\mathinner{.\,.}|T|)$ (Type 2) we use RMQs, similar to Section 3. Namely, each occurrence is reported in $\mathcal{O}(1)$ time. For the occurrences of $P$ with a starting position in $[0,q)$ and an ending position in $q,|T^{i}|)$ (Type 5), we retrieve the longest suffix of $P$ that is a prefix of $T[p+1\mathinner{.\,.}|T|)$ and the longest prefix of $P$ that is a suffix of $T[0\mathinner{.\,.}q)$, and trigger a prefix-suffix query. The total time to process the deletion is $\mathcal{O}(1)$ and for reporting is $\mathcal{O}(\textsf{Occ})$. After reporting the occurrences of $P$ in $T^{i}$, we do nothing (to revert the edit). The correctness follows from 1. We have arrived at Theorem 2.
@@
5 Pattern Matching with Ephemeral Substring Edits
In this section, we prove Theorem 3. We design a data structure for pattern matching with ephemeral substring insertions, deletions, or substitutions. First, we preprocess two strings $T$ and $P$ each of length at most $n$, over alphabet $\Sigma=[0,\sigma)$ with $\sigma=n^{\mathcal{O}(1)}$. Then, we allow any ephemeral sequence of edit operations in $T$. Before reverting the $i$th operation, we report all Occ occurrences of $P$ in $T^{i}$ in the optimal $\mathcal{O}(\textsf{Occ})$ worst-case time. This improves upon Theorem 1 for this setting.
@@ 5.1 Preprocessing
We start by constructing the data structure of Theorem 4 for prefix-suffix queries over $P$. This takes $\mathcal{O}(|P|)$ time and space. We also construct the String Matching Automaton (SMA) of the reverse of $P$. The SMA of a string $X$, denoted by $\textsf{SMA}(X)$, is the minimal deterministic automaton that accepts the language $\Sigma^{*}X$, where $X\in\Sigma^{*}$. It has $|X|+1=\mathcal{O}(|X|)$ states, one for every prefix of $X$, with $\varepsilon$ being the initial state and $X$ being the accepting state. The crucial property of SMA is that only the transitions that do not lead to the initial state are stored. There are at most $2|X|=\mathcal{O}(|X|)$ of them, regardless of the size of the alphabet [CHL07]. Moreover, the $\textsf{SMA}(X)$ can be constructed in $\mathcal{O}(|X|)$ time [CHL07]; its transitions can be accessed in $\mathcal{O}(1)$ time if stored using perfect hashing [BCF+23]. This preprocessing takes $\mathcal{O}(|P|)$ time and space.
Deletions are handled by Theorem 2; we thus shift our focus to preprocessing for insertions. We employ 2 (inspect Figure 8). We construct the suffix tree $\textsf{ST}(T\#P)$ of T#PT\#P, $\notin\Sigma$, and mark the nodes of the suffix tree that represent the suffixes of $P$. Every marked node representing $P[i\mathinner{.\,.}|P|)$ has a pointer to the state of the SMA representing the reverse of $P[i\mathinner{.\,.}|P|)$. Using a bottom-up traversal on the suffix tree and a stack to maintain the currently active marked nodes, we store a pointer from every node $v$ representing a suffix of $T$ to its deepest marked ancestor (i.e., the longest suffix of $P$ represented by an ancestor of node $v$). We also store the longest prefix of $P$ which is a suffix of the prefix $T[0\mathinner{.\,.}p]$ of $T$, for all $p$ (see Section 4.1). This preprocessing takes $\mathcal{O}(n)$ time and space.
@@
Figure 8: On the left is the suffix tree of $T\#P=\texttt{bababbbababb}\#\texttt{ababab}$. The label of edges leading to leaf nodes is truncated to avoid cluttering the figure. The nodes representing suffixes of $P$ are colored red. On the top right is the SMA of the reverse of $P$. Consider the query $p=4$ and $c=\texttt{a}$. The pointer from the leaf node labeled $p+1=5$ of the suffix tree takes us to its deepest colored ancestor labeled 1818; the corresponding state in the SMA is 1. From there, if we follow the a transition, it takes us forward to state 2, meaning that the longest suffix of $P$ preceded by $c=\texttt{a}$, represented by an ancestor of the leaf node labeled 55, is b, which yields ab.